OpenStudy (anonymous):

x: 2 4 6 9 11 y: 23 5 13 25 33 1. Explain why the data can be modeled by a linear equation. 2. Write an equation in point-slope form that relates y to x

OpenStudy (shamil98):

Not sure about 1 ( a bit tired atm) but i'll help with #2 Take the slope of two pairs of coordinates, let's use (2,23) and (4,5) calculate the slope and enter the coordinates of (x1,y1)-->(2,23) and m -->slope into y - y1 = m(x-x1) to get point-slope form

OpenStudy (anonymous):

Ok and is the slope -9?

OpenStudy (shamil98):

5-23/4-2 = -18/2 = -9 yes correct!

OpenStudy (anonymous):

Thank you random math genius!